OpenStudy (anonymous):

3(x-45)= 18x When I distribute the 45 to the 3, is it "-45"or "45"?

hartnn (hartnn):

3 (x-45) so 3 gets multipled by x and it get mutiplied by -45 then add both the terms

OpenStudy (anonymous):

Okay, so it would be: 3x - 135= 18x -3x -3x -------------------- -135/5x x=-27

hartnn (hartnn):

18-3 = ?

OpenStudy (anonymous):

Hmm this seems incorrect.

hartnn (hartnn):

yeah, 18x-3x = ... ?

hartnn (hartnn):

its not 5x

OpenStudy (anonymous):

Lol! Ohh opss.

OpenStudy (anonymous):

It's 15.

hartnn (hartnn):

yeah, try again..

OpenStudy (anonymous):

Okay, so my answer would be x=-9

hartnn (hartnn):

i get that too :)

OpenStudy (anonymous):

Thanks for all the help today! I appreciate it! Gbu!

hartnn (hartnn):

hey, most welcome ^_^ always happy to help :)